The electric field inside a 30-cm-long copper wire is 0.010 V/m.
What is the potential difference between the ends of the wire?
Answer #1
Concepts and reason

The concept needed to solve this problem is definition of electric field in terms of change in potential over a displacement.

Use the definition of electric field in terms of change in potential over a displacement and solve for required potential.

Fundamentals

The magnitude of electric field (E) is equal to amount of change in potential ( ΔV\Delta V ) over a displacement ( Δr\Delta r ) along the electric field direction.

E=ΔVΔrE = \frac{{\Delta V}}{{\Delta r}}

The expression for the electric field is,

E=ΔVΔrE = \frac{{\Delta V}}{{\Delta r}}

Rearrange the equation for ΔV\Delta V .

ΔV=E(Δr)\Delta V = E\left( {\Delta r} \right)

The final expression for potential difference between the ends of wire is,

ΔV=E(Δr)\Delta V = E\left( {\Delta r} \right)

Convert the unit of length of the wire from cm to m.

Δr=30cm(1m100cm)=0.30m\begin{array}{c}\\\Delta r = 30{\rm{ cm}}\left( {\frac{{1{\rm{ m}}}}{{100{\rm{ cm}}}}} \right)\\\\ = 0.30{\rm{ m}}\\\end{array}

Substitute 0.010 V/m for E and 0.30 m for Δr\Delta r in the equation ΔV=E(Δr)\Delta V = E\left( {\Delta r} \right) .

ΔV=(0.010V/m)(0.30m)=0.003V\begin{array}{c}\\\Delta V = \left( {0.010{\rm{ V/m}}} \right)\left( {0.30{\rm{ m}}} \right)\\\\ = {\rm{0}}{\rm{.003 V}}\\\end{array}

Ans:

The required of potential difference is 0.003 V.